OpenStudy (anonymous):

what is the measure of angle 1

OpenStudy (wolf1728):

You'll have to post a graphic.

OpenStudy (anonymous):

how do i do that

OpenStudy (wolf1728):

When you post, click on "Attach File"

OpenStudy (wolf1728):

You were successful

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do you know it

OpenStudy (wolf1728):

The triangle to the left of angle 1 and angle 2 is an equilateral triangle (all angles = 60 degrees). The bottom right angle is 90 degrees. The upper right angle must add up to 90 degrees. it consists of an equilateral triangle angle (60 degrees) plus angle 2 Therefore Angle 2 = 30 gegrees So, Angle 1 = 30 degrees.

OpenStudy (wolf1728):

Angle 1 = SIXTY Degrees

OpenStudy (anonymous):

i all ready put 30

OpenStudy (wolf1728):

Well, I was hurrying to finish that explanation

OpenStudy (anonymous):

thanks and i have more question

OpenStudy (wolf1728):

Okay - and make SURE I have the time to check my answer.

OpenStudy (anonymous):

is this the secound one

OpenStudy (wolf1728):

side BC

OpenStudy (anonymous):

so i was right

OpenStudy (wolf1728):

The rule is the shortest side is opposite the smallest angle

OpenStudy (wolf1728):

That drawing was made incorrectly. (Probably to throw off the folks that want to guess the answer from the drawing.) Angle A = 52 Angle B 62 However they drew angle A larger than Angle B

OpenStudy (anonymous):

is this never

OpenStudy (wolf1728):

Gee I'm not too sure about that (nothing to calculate) I'd say always (but not sure)

OpenStudy (wolf1728):

Okay, I thought about it and I'd say sometimes

OpenStudy (anonymous):

do you know this one

OpenStudy (wolf1728):

I'd say sometimes. Isosceles triangle 2 sides equal Equilateral 3 equal sides BUT I'd say Isosceles triangle can soemtimes have 3 equal sides

OpenStudy (anonymous):

i never understood this one

OpenStudy (wolf1728):

Okay the 2 sides marked with the line are equal. The base angles are equal. Therefore, 6x + 8 = 9x -25 33 = 3x x = 11 Left angle = 6*11 + 8 = 74 Right Angle = 9*11 - 25 = 74 BOTH base angles = 74 degrees

OpenStudy (anonymous):

thanks so much bc i never knew that one and do you know this one

OpenStudy (wolf1728):

I'd say angle ADE

OpenStudy (anonymous):

what about this one

OpenStudy (wolf1728):

All 3 angles add to 180 3rd angle = 180 -54 - 63 = 63 degrees

OpenStudy (anonymous):

I'm almost done

OpenStudy (wolf1728):

I'd have to say never based on this diagram.

OpenStudy (anonymous):

ya i think its never

OpenStudy (wolf1728):

Okay - anymore - cause I'll have to get going

OpenStudy (wolf1728):

Angle C

OpenStudy (wolf1728):

segment PD

OpenStudy (wolf1728):

Gotta go - glad to help you out

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@Princess_Rose